OpenStudy (anonymous):

find the inverse matrix of A=[-1 3] -4 -3

OpenStudy (anonymous):

For matrix \[ \left( \begin{array}{cc} a & b \\ c & d \end{array} \right) \] inverse is \[\frac{1}{ad-bc} \left( \begin{array}{cc} d & -b \\ -c & a \end{array} \right)\]